OpenStudy (anonymous):

4(x^2 - 9)^2 - (x + 3)^2/x^2 + 6x + 9

OpenStudy (anonymous):

Is the following the problem expression?\[\frac{4 \left(x^2-9\right)^2-(x+3)^2}{x^2+6 x+9} \]

OpenStudy (anonymous):

Yes @robtobey

OpenStudy (anonymous):

Multiply out the numerator.\[\frac{4 x^4-72 x^2+324-x^2+6 x+9}{x^2+6 x+9} \]\[\frac{4 x^4-72 x^2+324-x^2+6 x+9}{x^2+6 x+9}=\frac{4 x^4-72 x^2+324}{x^2+6 x+9}-\frac{x^2+6 x+9}{x^2+6 x+9}=\frac{4 x^4-72 x^2+324}{x^2+6 x+9}-1 \]\[\frac{4 x^4-72 x^2+324}{x^2+6 x+9}-1=\frac{\left(4 (x-3)^2 (x+3)^2\right)}{(x+3)^2}-1=\left(4 (x-3)^2 \right)-1=4 x^2-24 x+35 \]

OpenStudy (anonymous):

Rob the answer is (2x - 7)(2x - 5)

OpenStudy (anonymous):

I don't know how to derive it.

OpenStudy (anonymous):

\[4 x^2-24 x+35=(2 x-7) (2 x-5) \]

OpenStudy (anonymous):

I can't see what you type, Open study cut it off.

OpenStudy (anonymous):

I need to fix this problem before I can see people's comments. /sigh

OpenStudy (anonymous):

The following\[4 x^2-24 x+35 \]can be factored to give the "answer"\[(2 x-7) (2 x-5) \]Did the above come through?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But I want to see the method you did for solving the problem but I can't because the side is cut off. Do you mind retyping it but in a way so it doesn't get cut off?

OpenStudy (anonymous):

Hold on. I will attach a Mathematica PDF file.

OpenStudy (anonymous):

@robtobey Are you still there?

OpenStudy (anonymous):

Almost finished.

OpenStudy (anonymous):

A Mathematica solution is attached.

OpenStudy (anonymous):

You could also solve by division. Refer to the attachment.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome.